### v1

# Prompting in Practice: Investigating Software Developers’ Use of Generative AI Tools

Abstract.
The integration of generative artificial intelligence (GenAI) tools has fundamentally transformed software development. Although prompt engineering has emerged as a critical skill, existing research focuses primarily on individual techniques rather than software developers’ broader workflows. This study presents a systematic investigation of how software engineers integrate GenAI tools into their professional practice through a large-scale survey examining prompting strategies, conversation patterns, and reliability assessments across various software engineering tasks.
We surveyed 91 software engineers, including 72 active GenAI users, to understand AI usage patterns throughout the development process. Our ???? key findings show that while code generation is nearly universal, proficiency strongly correlates with using AI for more nuanced tasks such as debugging and code review, and that developers prefer iterative multi-turn conversations to single-shot prompting. Documentation tasks are perceived as most reliable, while complex code generation and debugging present sizable challenges. Our insights provide an empirical baseline of current developer practices, from simple code generation to deeper workflow integration, with actionable insights for future improvements.
1. Introduction
Generative Artificial Intelligence (GenAI) is rapidly transforming the way developers create and evolve software. With the help of GenAI tools such as GitHub Copilot (GitHub, [n. d.]), Windsurf (codeium, [n. d.]), and ChatGPT (OpenAI, [n. d.]), developers can now easily and quickly generate new code, review and modify existing codebases, identify and correct defects, and produce code documentation. Central to these capabilities is prompting: the practice of crafting natural language instructions to guide GenAI systems toward producing useful and accurate output. Prompting has emerged as a novel engineering practice, fundamentally different from traditional software development due to its reliance on large language models (LLMs), which are inherently probabilistic and non-deterministic  (Atil et al., 2025).
Previous research has primarily focused on cataloging prompting techniques (White et al., 2023; Schulhoff et al., 2025; Sahoo et al., 2025), which offers limited insight into how developers actually use prompting in practice: how they craft prompts for different types of software engineering (SE) tasks, how they adapt to unreliable or suboptimal output, and how GenAI is reshaping standard development activities. Previous studies have examined narrow use cases (e.g., refactoring (AlOmar et al., 2024)), specific populations (e.g., students (Alpizar-Chacon and Keuning, 2025; Choudhuri et al., 2025a)), and broader socio-organizational adoption factors (Li et al., 2024; Russo, 2024), but there remains a lack of systematic investigation into real-world prompting practices in the full range of SE tasks.
To address this gap, we conducted a comprehensive survey of 72 software developers, exploring how they use GenAI tools and engage in prompting across six common SE tasks: code generation, documentation, debugging, testing, refactoring, and review. Our study investigates various dimensions of GenAI usage, including the prompting and conversational strategies developers employ, how they assess and respond to the quality of GenAI outputs, and how prompting practices are integrated into their daily workflows. Together, these dimensions provide a holistic view of how prompting functions in modern software development and extend the current understanding of prompt engineering in practice.
We find that while nearly all developers who use GenAI apply it to code generation, higher perceived proficiency strongly correlates with engagement in more types of tasks, such as code refactoring and testing. Developers rely primarily on iterative conversations, with daily usage critical to achieving significant productivity improvements. Our reliability assessment reveals that developers trust GenAI most for documentation and least for complex code generation/modification tasks. These findings have important implications for developers seeking to improve productivity through effective prompting strategies, for tool builders designing more reliable and intuitive GenAI, and for educators developing curricula around the evolving skill set needed for effective GenAI collaboration.
This paper makes the following contributions: (1) a comprehensive characterization of how GenAI tools are used in various SE tasks; (2) empirical insights into developers’ prompting strategies, reliability perceptions, and common failure types; and (3) actionable implications for improving human-AI interaction in SE.
2. Study Methodology
The goal of our study is to investigate how developers use prompt engineering and GenAI tools for development tasks and what challenges they face during this use. The context consists of software developers with experience using GenAI tools for SE.
Our study addresses the following research questions (RQs):
  1. • RQ1:
    How do software developers integrate GenAI tools into their development workflows?
  2. • RQ2:
    What prompting techniques and conversation strategies do developers use when interacting with GenAI?
  3. • RQ3:
    How reliable do developers perceive GenAI to be for various SE tasks and what are common issues they encounter?
Below, we detail our methodology (Figure 1), which aligns with our prior studies (Stalnaker et al., 2024a, b; Wintersgill et al., 2024; Zappin et al., 2025) and was approved by our institution’s ethics review board, including the questionnaire, participant identification procedure, and data collection/distribution.
Figure 1. Study Methodology (image credits at (Otten et al., 2025))
2.1. Survey Design
We designed our survey questionnaire considering existing literature on prompt engineering and GenAI tool use (White et al., 2023; Schulhoff et al., 2025), following general guidelines (Groves et al., 2009) and SE-specific guidelines for survey design (Pfleeger and Kitchenham, 2001; Kitchenham and Pfleeger, 2002a, b, c, d, 2003). Through the survey, our goal was to build an understanding of GenAI integration across six core SE tasks: code generation, refactoring, testing, debugging, documentation, and code review. Although other SE tasks exist, we opted to include only these six in an attempt to balance survey length and appropriate task coverage. We also implemented branching logic in the survey, so that participants were only shown questions relevant to their experiences. To ensure data quality, we also included an initial question to screen out respondents with no experience using prompt-based GenAI tools. We carefully formulated survey questions, ensuring that they were written clearly to avoid confusing or biasing respondents. We also conducted a pilot study with graduate students from our research lab who regularly use GenAI tools. Based on their feedback, we further refined our questionnaire, improving questions that were unclear or confusing.
The survey was organized into five sections. Section 1 collected demographic information. Sec. 2 asked questions related to the GenAI experience of the participants, the GenAI tools they regularly use, and their GenAI development workflows. Sec. 3 and 4 investigated the use and familiarity of participants with prompting techniques and conversation strategies. The final section explored the perceived reliability of GenAI tools and the issues that developers face. Questions were presented in a variety of formats, including multiple-selection, Likert scales, and open text fields. The full text and logic of our survey can be found in our replication package (Otten et al., 2025).
2.2. Participant Recruitment
Our recruitment strategy targeted developers with diverse backgrounds and experience levels. For survey distribution, we mainly relied on our professional networks and snowball sampling (e.g., participants sharing the survey within their networks) by directly contacting potential participants or posting participation calls in social media platforms (e.g., LinkedIn and X). To further broaden our sample and include participants at earlier career stages, we distributed the survey to relevant organizations at our university, including the Association for Computing Machinery (ACM) chapter. Participants were not compensated monetarily or otherwise.
2.3. Data Collection and Analysis
Survey responses were collected using Qualtrics (qua, nd) between 27 May and 7 July 2025. To analyze the data, we employed a mixed method approach, combining quantitative and qualitative techniques.
For multiple choice and Likert scale questions, we calculated data distributions and descriptive statistics to identify trends and patterns. Were appropriate, to facilitate comparison, we assigned ordinal Likert scale responses (e.g., ”Often” to ”Never”) to integer values (e.g., 1 to 4). This allowed us to compute mean scores to represent concepts such as the popularity of a prompting strategy or the frequency of an issue using GenAI tools.
The responses to open-ended text responses were analyzed using a qualitative coding approach (Spencer, 2009). Three annotators (i.e., authors) independently analyzed open-ended text responses and performed open-coding by assigning one or more codes to each response using a shared codebook (i.e., a spreadsheet). After the initial coding phase, the annotators met to collaboratively review, discuss and reconcile the coding decisions. Through this process, the annotators resolved any discrepancies, which included differing interpretations of potentially ambiguous responses, applying the same codes in different contexts, and using broader codes rather than more specific ones. Ultimately, this process resulted in a final set of codes, definitions, and labels for each response, which are found in our replication package (Otten et al., 2025). Since codes were developed iteratively through an inductive process and multiple codes could be assigned to each response, we did not base our analysis on inter-annotator agreements. We followed best open-coding practices (Spencer, 2009) and used annotator discussions to ensure the reliability of the results.
GenAI Proficiency Code Gen. Refactoring Testing Debugging Documentation Code Review Avg. # of tasks
Minimum Proficiency 1/1 (100%) 0/1 (0%) 0/1 (0%) 0/1 (0%) 0/1 (0%) 0/1 (0%) 1.00
Somewhat Proficient 11/14 (78.5%) 3/14 (21.4%) 6/14 (42.9%) 4/14 (30.8%) 4/14 (30.8%) 3/14 (23.1%) 2.21
Proficient 32/32 (100%) 11/32 (34.4%) 9/32 (28.1%) 16/32 (50%) 14/32 (43.8%) 8/32 (25.0%) 2.81
Very Proficient 15/18 (83.3%) 6/18 (33.3%) 8/18 (41.2%) 8/18 (41.2%) 11/18 (58.8%) 8/18 (41.2%) 3.11
Maximally Proficient 7/7 (100%) 3/7 (42.9%) 5/7 (71.4%) 4/7 (57.1%) 5/7 (71.4%) 4/7 (57.1%) 4.00
Total 66/72 (91.7%) 23/72 (31.9%) 28/72 (38.9%) 32/72 (44.4%) 34/72 (47.2%) 23/72 (31.9%) 2.86
Table 1. Proportion of respondents in each self-reported proficiency level who indicated GenAI use for each SE task
2.4. Participant Demographics
Our survey attracted 91 respondents who completed the survey, of whom 72 (79%) indicated having experience with GenAI tools and prompting. We answer our RQs based on these 72 responses.
The 72 participants represent a diverse, global sample of software professionals, with reported experience ranging from one to 40 years (mean = 10.2, median = 10). They held a variety of roles, including researcher (64%), programmer (60%), technical lead (28%), and software architect (28%). Respondents could select multiple roles, and overlaps were common: for example, while 46 participants identified as researchers (the most frequently selected role) 31 of them also reported other roles, notably programmer, architect, technical lead, and/or tester. For comparative analysis, we grouped professional experience into four categories: 24 juniors (1–5 years), 22 intermediates (6–10 years), 19 seniors (11–20 years), and 7 veterans (21+ years).
Participants reported having developed software across 43 distinct domains, with the most common being Healthcare (22%), Research (17.6%), Banking (15.4%), Software Development (12.4%), and Communication (10.1%). When asked to choose from a predefined list of eight system types, they most frequently indicated experience with web applications (63.9%), desktop applications (33.3%), libraries/frameworks (31.2%), and AI-intensive systems (30.6%). Respondents were located in 16 countries, with the majority residing in the United States (54%), followed by Canada (11%) and the United Kingdom (8%). The most commonly used programming languages were Python (83%), Java (61%), and JavaScript (39%).
3. Results
3.1. RQ1: GenAI Usage Patterns
3.1.1. GenAI Use for SE Tasks and Proficiency
Table 1 shows that the vast majority of respondents (91.7%) regularly use GenAI for code generation. Although this is somewhat expected, we found that debugging (47.2%) and code documentation (44.4%) emerged as next most common tasks. The remaining tasks, testing (38.9%, refactoring (31.9%), and code review (31.9%) were less frequently reported. Overall, 89% of the respondents reported using GenAI for four or fewer tasks, with two (30.6%) or three (23.6%) tasks being most common, suggesting a relatively focused use of GenAI.
When analyzing the 5-Likert point responses of self-reported proficiency in using GenAI, we found that the majority of participants (57 of 72) considered themselves proficient or very/maximally proficient (see Table 1). (Since a single veteran developer rated themselves as having “minimum proficiency”, we exclude this data point from future comparisons across proficiency levels.)
We also observed that proficiency correlates with the number of SE tasks for which participants use GenAI.(Spearman’s $\rho$ = 0.325). Among those who reported using GenAI for four or more SE tasks, most (18 of 22, 82%) identified as very or maximally proficient. By contrast, single-task users reported being somewhat proficient or proficient (7 of 11, 64%). These results suggest that task breadth and GenAI proficiency may mutually reinforce one another.
When examining proficiency by software development experience, we found that junior and intermediate developers considered themselves more proficient in using GenAI that seniors and veterans. Specifically, 22 of 24 juniors and 19 of 22 intermediates reported being proficient or very/maximally proficient, compared to only 12 of 19 seniors and 4 of 7 veterans.
Finding 1: A moderate correlation between perceived GenAI proficiency and GenAI use across six SE tasks suggests that that greater proficiency is associated with broader usage, and vice versa. Junior and intermediate developers tend to rate themselves as more proficient in using GenAI compared to senior and veteran developers.
3.1.2. GenAI Usage Frequency and Tool Adoption
Over 70% of respondents reported engaging with GenAI tools once or multiple times per day. Daily usage increases with proficiency: 35.7% of somewhat proficient users (5) reported daily GenAI use, as did 78.1% of proficient users (25), 83.3% of very proficient users (15), and 100% of maximally proficient users (7).
Figure 2. Productivity and task breadth by usage frequency
Respondents who reported using GenAI more frequently tended to also report using GenAI for more SE task and experiencing productivity gains (Figure 2). Daily users reported using GenAI for 3.12 tasks on average, compared to 2.42 tasks for weekly user (an increase of 29%). Reported usage frequency was moderately correlated with the breadth of GenAI use across tasks (Spearman’s $\rho$ = 0.348). Compared to non-daily users, daily GenAI users were also substantially more likely to state that GenAI significantly increased their productivity, with only one respondent who used GenAI tools non-daily reported such substantial gains. Taken together, these results suggest that sporadic GenAI use is insufficient to realize meaning productivity gains and that users who perceive higher productivity gains are, in turn, more likely to use such tools often.
Finding 2: Most respondents (69.3%) reported using GenAI for SE tasks once or multiple times per day. In contrast, only 35.7% of those who considered themselves somewhat proficient reported daily GenAI usage, suggesting a link between consistent use and perceived proficiency. Respondents who used GenAI more frequently also reported applying it across a larger number of SE tasks.
Analyzing tool adoption from a given predefined list of options (with an open-ended ”Other(s)” option), OpenAI’s ChatGPT led reported tool adoption (83.1%), followed by GitHub Copilot (43.7%), Google’s Gemini (35.2%), and Anthropic’s Claude (32.4%). Open-weight models such as Deepseek-R1 (12.7%) and the Llama family of models (7%) showed more modest adoption. The use of AI-integrated development environments (besides GitHub Copilot) was reported by 19.7% of the participants. Examples include Cursor (6), Windsurf (1), and JetBrains AI Assistant (1).
Finding 3: The vast majority of developers (83.1%) reported using ChatGPT for SE tasks, which was nearly 2× higher than the next most popular tool (GitHub Copilot with 43.7%), suggesting that it has established itself as the leading platform for AI-assisted software development.
When interacting with GenAI tools for SE tasks, an equivalent number of participants reported using tools with a web interface (46, 63.9%) as an IDE-integrated solution (45, 62.5%). Other GenAI integrations such as GenAI APIs (14, 19.4%) and command line tools showed much lower adoption (13, 18.3%). Despite having used GenAI for development tasks, five respondents (6.9%) indicated they had not been incorporated it into their workflows.
A majority of developers (39, 54.2%) reported integrating GenAI into their workflows in two or more ways, while 28 (38.9%) reported a single integration, most commonly IDE-based (13) or web-based (12). Among participants using two integration types, 66.7% (20) reported using both web-based and IDE-integration. These findings suggest that while command-line and API-based GenAI integrations are available, they are rarely used in isolation by developers. Table 2 summarizes the # of participants across GenAI tool integrations. (No ”Other” integrations were reported by participants.)
GenAI Integration # of Users Percentage
No integration 5 6.9%
Single integration 28 38.9%
Dual integration 30 41.7%
Triple integration 6 8.3%
Four integrations 3 4.2%
Table 2. Distribution of participants by number of GenAI workflow integrations (Web, IDE, Command line, and API)
Respondents that reported daily GenAI usage also demonstrated a higher number of workflow integration types when compared with weekly users. 60.9% (31/51) reported used two or more integration methods, compared to 25% (3/12) for weekly users.
Finding 4: Developers reported no preference between incorporating web-based or IDE-integrated GenAI tools into their development workflows, but did overwhelmingly employ one or both over command-line and API-based tools. Developers who reported using GenAI tools multiple times a day were also more likely to incorporate GenAI tools into their workflows in multiple ways (e.g., Web-based, IDE-integration, etc.). This suggests that frequent GenAI users may develop more tailored or complex workflows.
3.1.3. Changes in Development Approaches with GenAI
When asked if GenAI tools had changed their approach to software development, the majority of respondents (55/71, 77.5%) indicated that they had experienced at least some change, while 7.0% were unsure and 15.5% reported no change. We asked respondents who reported experiencing changes to their development approach to elaborate on those changes. Nine participants reported using GenAI tools for tasks that previously required internet searches (e.g., finding library documentation or searching Stack Overflow for code examples), and four reported using GenAI to assist in information retrieval tasks within their code base. Six indicated that GenAI can provide a starting point for projects, such as helping with ideation (3), initial project design (5), or writing boilerplate code (6). Five respondents noted that GenAI allowed them to shift development focus from repetitive coding tasks to higher-level tasks such as architecture design, and four respondents claimed that GenAI tools allowed them to write more documentation. Still others indicated that GenAI helps produce cleaner, more organized code (3), reduces the focus and memory required to complete coding tasks (2), allows for development in unfamiliar programming languages (2), and can assist in project configuration (2).
GenAI tool adoption has created new processes in developer workflows. For example, six respondents spend more time reviewing and verifying AI-produced code. Four focus more on writing good prompts, while three have incorporated interactive prompting processes into their workflows.
Finding 5: Developers reported that GenAI has transformed their development workflows in diverse ways, ranging from replacing traditional search practices to supporting project design, reducing repetitive effort, and broadening programming capabilities, ultimately enabling a shift toward higher-level, more strategic software engineering tasks. In addition, participants noted that GenAI introduced new practices in their development workflow, such as validating AI-generated outputs and employing various prompt engineering strategies.
3.2. RQ2: Prompting and Conversation Strategies
This section examines the specific prompting techniques and conversation strategies that software developers use when interacting with GenAI tools. We first analyze participants’ familiarity with prompting strategies, then investigate what contextual information developers include when prompting GenAI for different SE tasks, what conversation strategies they use, and how developers react to inadequate output from GenAI tools.
3.2.1. Prompting Technique Familiarity
Participants were asked about three prompting techniques identified by White et al. (Meta Language Creation, Output Automator, and Persona(White et al., 2023) in addition to Meta-Prompting and Few-Shot Learning (Schulhoff et al., 2025). We also included the Output Style and Condition Check techniques, which we did not find adequately represented in existing work. Definitions for these techniques were provided to participants and can be found in Table 3.
Prompting Technique Technique Description
Condition Check Produce certain type(s) of output when specific condition(s) are met
Few-Shot Learning Providing examples to guide generation
Meta Language Creation Adjusts/Adds to the semantics of words, phrases, or symbols
Meta-Prompting Having AI suggest better prompts for specific tasks
Output Automator Generating scripts to implement solutions
Output Style Make the AI’s output follow a particular format or style
Persona Ask AI to complete a task while acting as a certain role or character
Table 3. Prompting technique descriptions
Figure 3. Prompting technique familiarity distribution
Our analysis of prompting technique familiarity reveals a clear hierarchy of adoption, as shown in Figure 3. We consider participants were familiar with a technique if they knew or used the technique (occasionally or regularly). Respondents reported overwhelming familiarity with the aforementioned prompting strategies. Nearly all respondents were familiar with the Few-Shot Learning (93%), Output Style (92%), and Persona (92%) techniques. The Meta Language Creation technique were the least familiar to participants (74%). Despite the high familiarity, reported respondent usages (either occasionally or regularly) were markedly lower. The Output Style (82%), Few-Shot Learning (75%), Persona (68%), and Condition Check (65%) strategies had the highest reported usage rates among respondents who were familiar with the respective strategies. Roughly half of respondents familiar with the Output Automator (51%) and Meta-Prompting (49%) strategies reported using them. The Meta Language Creation strategy had the lowest adoption rate (36%).
The Output Style (54%) and Persona (44%) strategies were most likely to be used regularly, among participants who used them. Despite being least popular overall, 42% of respondents who used the Meta Language Creation strategy do so regularly. Respondents most often reported using the Output Automator (75%), Condition Check (68%), and Meta-Prompting (68%) strategies only occasionally.
Finding 6: Respondents were highly familiar with all seven prompting strategies, with 74%–94% reporting awareness. However, actual usage lagged behind, with only 36%–82% of those aware indicating that they used the techniques.
Proficiency Familiar % Use % Occ. Use % Reg. Use %
Somewhat 74.5 52.1 89.5 10.5
Proficient 83.9 60.1 62.8 37.2
Very 89.7 69.0 51.3 48.7
Maximally 95.9 68.1 34.4 65.6
Experience Familiar % Use % Occ. Use % Reg. Use %
Junior 85.1 58.0 67.5 32.5
Intermediate 87.7 63.7 45.4 54.7
Senior 86.5 62.6 66.7 33.3
Veteran 57.1 71.4 65.0 35.0
Table 4. Prompting technique familiarity and use by GenAI proficiency and SE experience categories
We found that none of respondents’ self-reported GenAI proficiency, years of development experience, or GenAI proficiency and experience impacted relative familiarity ratings of prompting techniques (e.g., across proficiency groups, Output Style always had a higher familiarity than Meta Language Creation).
Table 4 reports the percentage of respondents who are familiar with, and use, the seven prompting techniques across GenAI proficiency and SE experience categories. Technique usage rates are further reported by occasional and regular use. The table reports total percentages by aggregating response counts across the techniques. The results show a clear pattern: both technique familiarity and usage (including occasional and regular use) consistently increase with self-reported proficiency. In other words, participants who considered themselves more proficient were also more likely to use prompting techniques more frequently.
In contrast, analysis by developer experience reveals a different trend. With the exception of veterans, all experience groups show similar familiarity and usage rates. Veteran are the least familiar with prompting techniques, but among those who are familiar, they are the most likely to use them. Analyzing occasional versus regular technique usage, developers across all experience levels show consistent rates, except for intermediate developers, who reported over 60% increase in regular usage compared with their peers. This suggests that intermediate developers may be at a career stage where they are both open to adapting their workflows and sufficiently experienced to integrate GenAI effectively.
Finding 7: Familiarity and usage of prompting techniques increase with GenAI proficiency rather than years of experience. Veteran developers are the least familiar with prompting techniques, yet the most likely to use the ones they know. While most experience groups use prompting at similar rates, intermediate developers stand out as especially active adopters, suggesting they are mid-career professionals that are uniquely positioned to integrate GenAI into their workflows.
3.2.2. Prompt Information Inclusion
We describe the prompt information developers use for each of the six SE tasks we studied.
When prompting a GenAI tool for code generation, 76% of respondents (66) indicated that they often include example inputs and expected outputs. Respondents also indicated including instructions for which libraries the model should use (53%), a description of the existing codebase (45%), implementation alternatives to either consider or avoid (41%), error handling expectations (39%), coding style guidelines (32%), deployment constraints (32%), and performance/optimization requirements (27%).
When using GenAI tools for refactoring tasks, respondents (23) most often included the original code with explanatory comments (78%), a description of the existing issues or refactoring goals (65%), and examples of similar code that they found to be well structured (61%). Others opted to include architectural constraints and style guidelines (48%), a specific refactoring pattern to apply (48%), or code that should remain unchanged during the refactoring (43%). The least common information to be supplied in respondent prompts included unit tests (22%) and performance requirements (17%).
Of the 28 respondents that used GenAI for testing tasks, 82% prompted the model with the code to be tested along with specifications. Respondents also included example test cases (46%), testing preferences (43%), information on edge cases or expected behaviors (39%), existing test suite examples (39%), and mocking instructions (36%) in their prompts. Least likely to be included were test coverage goals (25%), environmental requirements (25%), and performance or resource constraints (14%).
When using GenAI for debugging tasks, nearly all respondents (94%) provide the model with full error messages or stack traces. Similarly, 81% often included related logs or console output. Less common context given to the model included environmental details (50%), previously attempted solutions (44%), steps-to-reproduce (41%), system architecture information (31%), and visual evidence of issue manifestation (e.g., screenshots) (25%). Only two respondents (6%) included version control history.
For documentation tasks, respondents (34) reported that they often provided the model with the code snippet along with a functionality explanation (71%). Some also provided examples highlighting the desired documentation style (53%), usage examples/scenarios (41%), context about the target audience (38%), existing documentation in need of updating (38%), format requirements (35%), guidance on domain-specific terminology (26%), or information on project standards and conventions (15%).
Lastly, when using GenAI for code review, respondents (23) predominantly provided the model with the code to be review with some additional context (78%). Beyond this, they included project coding standards (43%), a description of the intended feature or purpose (35%), or performance expectations (30%), security/compliance requirements (22%), warnings about common pitfalls in the domain or codebase (22%), the code of related components (17%), or feedback from previous reviews (17%).
3.2.3. Conversation Strategy Adoption
We examined how developers structured their conversations with GenAI tools. Participants rated the use frequency of nine conversation strategies we defined (seen in Table 5) on a four-point Likert scale from “Never” to “Often.” This analysis reveals which strategies survey developers prefer.
Conversation Strategy Description
Comparative analysis Having the AI generate multiple solutions for comparison
Context building Adding more context as the conversation develops
Exploratory dialogue AI explores multiple solutions before implementing one
Feedback loop Providing feedback on AI responses to guide future outputs
Incremental refinement Iterative refinement on basic request based on AI responses
Multi-part problem solving Dividing the task into sub-problems addressed in sequence
Single comprehensive prompt One prompt with all requirements, context, and constraints
Step-by-step guidance Breaking problem down into sequential steps for AI to solve
Template-based approach Applying prompt templates adapted for task
Table 5. Conversation strategy descriptions
Figure 4 reveals that the most frequently used strategy was Incremental refinement (incrementally refining the same prompt) followed closely by Feedback loop (using potentially different prompts to achieve the desired result). These strategies reflect developers’ preference for multi-turn conversations where they can guide and refine GenAI responses progressively. However, Single comprehensive prompt is the third most popular and does not reflect this iterative workflow. The least popular strategies were Comparative analysis, Template-based approach, and Exploratory dialogue. Despite their lower adoption rates, these strategies still showed meaningful usage patterns, with at least 40% of developers employing them “Sometimes” or “Often”.
Figure 4. Conversation structure usage frequency
Figure 5. Error handling strategy usage frequency
3.2.4. Error Handling Strategies
We asked participants to indicate which error handling strategies they use when they deem GenAI responses as inadequate. Participants were provided with nine strategies and could select none or any number of strategies, including none. Developers converge overwhelmingly on two complementary strategies (Figure 5): providing additional context and pointing out specific issues, with 79% of frequent users employing both approaches. Figure 5 shows that reformulating requests, adding constraints, including example solutions, breaking down complex problems, and providing feedback on partial solutions show more moderate adoption, while requesting alternatives or explanations remain far less popular. However, the 13 respondents who often use this least popular strategy are likely to use every other strategy: none chose “Never” in response to any option. There is a similar trend for the 15 respondents who frequently request alternatives: only three selected “Never” for requesting explanations and not for any other strategy.
Finding 8: Developers who reported utilizing less popular error handling strategies typically also used all other provided strategies at least rarely. This suggests that strategies are used as complements to others.
The remainder of this section reveals connections between conversation structuring and error handling approaches, focusing exclusively on users who responded “Often” to both types of strategies. 80% of developers who frequently use Incremental refinement also provide additional context when errors occur. Similarly, 83% of users who provide feedback on partial solutions frequently utilize Feedback loop. Those who use step-by-step guidance often break down complex problems when handling errors (50%). These correlations show consistent developer strategies, which could inform future tool design: for example, a tool might activate additional context provision settings if a developer is aware of their preference for an iterative strategy, or adjust a prompting agent’s writing style depending on the user.
Two distinct conversation preferences emerge: Comprehensive users who rely on single detailed prompts (29/72) versus Iterative users who employ multi-turn refinement (48/72), with an overlap of 19 respondents.
Finding 9: Developers consistently reported combining related conversation structuring and error handling strategies, demonstrating preferences for two distinct GenAI interaction styles. Comprehensive prompters try to include as much detail in prompts as possible, while Iterative prompters start conversations planning to add additional context as needed while adapting to LLM outputs, with the Iterative style being significantly more popular.
3.2.5. Exchange Count
We asked respondents to report the average number of exchanges (back-and-forth interactions) needed to complete SE tasks with GenAI tools. Survey options included predefined ranges (1, 2-3, 4-6, 7-10, 10+, and Unsure). No participants reported consistently completing tasks in a single exchange, suggesting that current GenAI systems require iterative interaction regardless of user expertise or strategy choice. Self-reported GenAI skill levels show a positive correlation with exchange count, as seen in  Table 6. Maximally proficient users require substantially more exchanges than other skill levels, having by far the highest proportion of “10+” selections. Longer conversations also correlate with productivity gains: participants using 10+ exchanges were the only group to report exclusively positive productivity impacts, with 6/7 indicating “significant” improvements.
Finding 10: No developers indicated that they could consistently complete tasks with GenAI in a single exchange, while extended conversations with GenAI tools correlate with higher productivity and perceived skill.
GenAI Skill Level 2-3 4-6 7-10 10+ Unsure
Somewhat Proficient 3 7 0 0 4
Proficient 15 11 3 3 0
Very Proficient 5 9 1 1 2
Maximally Proficient 1 3 0 3 0
Total 24 30 4 7 6
Table 6. Exchange count distribution by self-reported skill
3.3. RQ3: GenAI Reliability and Common Issues
This section examines which SE subtasks developers perform using GenAI tools, their perceptions of these tools’ reliability across SE scenarios, and how often they encounter various issues. Our results highlight where GenAI tools excel and struggle.
3.3.1. Task Engagement Patterns
For each SE task, we analyzed which specific subtasks respondents reported using GenAI for (Table 7). The most common subtask for code generation was implementing new code components (52), followed by setting up structural code (34) and designing architectures (30). This pattern reflects a clear preference for using GenAI to assist with sub-tasks related to implementation rather than design/analysis. A similar implementation-focus emerges for some of the other SE tasks. For example, for testing, respondents showed a pronounced 6:1 preference for test creation/implementation (18) over test strategy/planning; and for documentation, a nearly 2:1 preference for code-level documentation (24) over project-level documentation (13). A few additional GenAI sub-tasks were not included in our predefined options, including reverse engineering, background research, and input generation for tests, each highlighted by one respondent.
Finding 11: For code generation, documentation, and testing, respondents generally reported using GenAI for implementation-focused (lower-level) sub-tasks rather than for design and analysis (higher-level) sub-tasks.
Table 7. Most frequently selected sub-tasks for each SE task. ‘I’ = Implementation focus (lower abstraction level)
‘D’ = Design / Review focus (higher abstraction level)
SE Task Sub-task Count
Code Generation Impl. new code components (I) 52
Setting up structural code (I) 34
Designing software architectures (D) 30
Documentation Code-level documentation (I) 24
Project-level documentation (D) 13
Testing Test creation and implementation (I) 18
Test strategy and planning (D) 3
Debugging Bug fixing (I) 26
Issue diagnosis (D) 22
Refactoring Optimizing existing code (I) 14
Standardizing and cleaning code (D) 12
Review Funct./perf. assessment (I) 10
Code quality evaluation (D) 7
Figure 6. Reliability perceptions of GenAI across tasks
3.3.2. GenAI Perceived Reliability
To assess developers’ trust in GenAI tools, we analyzed reported reliability ratings for 12 scenarios spanning all six studied SE tasks. Participants rated each scenario on a 4-point scale from “Very unreliable” to “Very reliable,” with “Unsure/Haven’t tried” responses excluded from our analysis.
Figure 6 shows the perceived reliability of using GenAI for various tasks. Respondents reported that documentation tasks were the most reliable to complete with GenAI. In particular, maintaining documentation accuracy received the highest avg. rating (3.06, n=32), followed by documenting implementation rationale (2.77, n=31). Respondents showed moderate confidence in using GenAI for testing and debugging tasks, including test case design and coverage (2.64, n=25) and root cause analysis of existing errors (2.57, n=30). The tasks reported as least reliable by developers often involved complex analysis or reasoning. These included the assessment of non-functional requirements (2.14, n=14), the modernizing of outdated code patterns (2.32, n=22), and the implementation of complex algorithms (2.35, n=60). In particular, GenAI tools may struggle with non-functional requirements since many of these require a deep understanding of domain knowledge from a complex field, such as security or license compliance.
Finding 12: Respondents perceived GenAI to be less reliable for software tasks that required complex reasoning, non-trivial analysis, and domain knowledge.
Experience Very- Some- Some+ Very+ Unsure Score
Junior 7.1% 22.7% 44.8% 12.8% 12.7% 2.73
Intermed. 12.2% 37.4% 36.0% 4.4% 10.1% 2.36
Senior 16.2% 35.3% 34.7% 2.6% 11.2% 2.27
Veteran 7.1% 19.1% 47.6% 17.3% 8.9% 2.82
Proficiency Very- Some- Some+ Very+ Unsure Score
Some. Prof. 29.5% 30.4% 27.4% 0% 12.8% 1.98
Proficient 6.2% 33.1% 42.8% 11.3% 6.6% 2.63
Very Prof. 8.4% 26.9% 37.7% 10.2% 16.7% 2.60
Max. Prof. 4.8% 29.1% 46.7% 4.1% 15.5% 2.59
Table 8. GenAI reliability results across GenAI proficiency and experience groups. (”Reliable” and ”Unreliable” shortened to ”+” and ”-”)
GenAI reliability results across respondents’ experience and self-reported GenAI proficiency groups can be seen in Table 8. Participants were presented with subsets of the 12 reliability options from Figure 6, which varied based on the SE tasks they indicated using GenAI for. To account for this variation of presented options, we calculated proportions for each participant across all of their reliability responses, then calculated a reliability score as the mean of individual reliability ratings across each GenAI proficiency and experience category. Overall reliability scores were calculated by multiplying integer values for the Likert-scale options (“Very unreliable” = 1, “Very reliable” = 4) by the calculated proportion of a given category, then dividing the % of participants who gave a rating other than “Unsure.” Overall reliability proportions and scores are reported in Table 8.
Despite nearly equivalent totals for Proficient, Very Proficient, and Maximally Proficient users, Somewhat Proficient users have the lowest at 1.98, largely due to an outlying quantity of “Very Unreliable” selections. This is expected, since it is unlikely that respondents would assign themselves low proficiency if they were able to obtain reliable AI responses. The most proficient users were unlikely to classify GenAI as “Very Unreliable” or “Very Reliable” for any task, suggesting that they have at least occasionally witnessed AI fail in optimal scenarios and succeed in sub-optimal ones.
In contrast, our analysis by experience level shows no clear patterns. The most- and least-experienced developers were more likely to find GenAI reliable than those in the middle. This may be correlated with Juniors’ and Veterans’ proportionally lower familiarity with prompting techniques, but that does not explain why Intermediate developers consider AI more reliable than Senior developers.
Finding 13: Senior developers (with 11-20 years of development experience) find GenAI least reliable on average, while veterans (with 21+ years) find it most reliable on average.
Figure 7. GenAI issue frequency distribution
3.3.3. Issue Frequency
Our final reliability analysis examined how frequently developers encountered specific issues (given to them) when using GenAI tools across different SE contexts. Participants rated twelve task-specific issues on a five-point scale from “Always” (1) to “Never” (5), where lower mean scores indicate more frequent issue occurrence. Results are visualized in Figure 7.
Debugging emerges as the most problematic SE task, with suggested bug fixes missing the root cause having the highest issue frequency (2.34, n=32). This issue received four “Always” responses (the highest count for any absolute rating across all issues, despite only appearing for those who selected the third-most popular task), indicating that GenAI’s diagnostic capabilities remain a critical hurdle. Refactoring and code generation follow in severity, with unintended changes to behavior (2.52, n=23) and correctness/compatibility issues in code (2.58, n=66) representing significant challenges. Documentation and code review showed the lowest reported issue frequencies, with developers noticing inaccurate or incomplete technical documentation (3.06, n=34) and false positives and hallucinations (3.00, n=23) least often. This aligns with our previous findings about the lower complexity of these tasks.
Finding 14: Respondents reported experiencing the most issues when using GenAI for debugging, particularly when it would miss the root cause of the problem. Code generation and refactoring were also noted pain points.
Table 9. GenAI issue frequency ranking across SE tasks
Rank Issue Mean n
1 Suggested bug fixes miss root causes 2.34 32
2 Unintended behavior changes 2.52 23
3 Code correctness and compatibility issues 2.58 66
4 Incomplete code transformations 2.61 23
5 Fixes that introduce new bugs 2.63 32
6 Codebase consistency issues 2.70 66
7 Superficial explanations of functionality 2.76 34
8 Low-value tests with inadequate coverage 2.82 28
9 Superficial code analysis 2.87 23
10 Tests that miss actual requirements 2.89 28
11 False positives and hallucinations 3.00 23
12 Inaccurate or incomplete documentation 3.06 34
4. Discussion
4.1. Developing Proficiency with Generative AI
Our results show a positive relationship between a developer’s perceived GenAI proficiency and the number of tasks they complete with GenAI (Table 1), suggesting that task diversity and GenAI proficiency may mutually reinforce each other. This could potentially reflect how confidence encourages developers to explore GenAI usage in new contexts, or could also demonstrate that using GenAI on a wide range of tasks increases proficiency.
Users’ prompt construction strategies vary significantly among tasks. Respondents showed a high consensus in the information they reported providing to models when completing debugging tasks, with 94% including full error messages/stack traces. By contrast, there was little agreement on the information included for documentation tasks, despite those being the second most common SE tasks completed with GenAI assistance: only 71% (24/34) of respondents reported providing the model with the code snippet and a functionality explanation. Code Generation has the next lowest with 76% of respondents including example inputs and expected outputs. All tasks show some convergence of information inclusion strategies, and debugging and testing (which have the highest agreement) explicitly involve work on existing code, making that necessary to share in most instances. However, Documentation also involves work on existing code in many cases, suggesting that consensus differences have no clear pattern.
One counterintuitive finding is that maximally proficient users require substantially more exchanges to complete tasks than less skilled users (Table 6), which seems to indicate that higher skill correlates with inefficiency. Another interpretation is that the strategies employed by these high-skill users are not inefficient, but naturally result in more prompts by enabling users to tackle more ambitious problems. Notably, longer conversations correlate with productivity gains: participants using 10+ exchanges were the only group to report exclusively positive productivity impacts, with 6/7 indicating “significant” improvements. This suggests that sustained engagement with GenAI tools enables more meaningful outcomes, rather than expecting immediate solutions.
4.2. Limitations and Practical Implications
Debugging does not have the lowest overall reliability perception, but it is prominent as the task with the highest issue frequency (#1 and #5 on Table 9), while other functional coding problems (refactoring and generation) fill the top 50%. This suggests consistent limitations in GenAI’s ability to reason about code behavior and the effects of modifications, compared to natural language tasks such as documentation and review.
While these results provide some direction for the critical problems GenAI engineers should focus on solving, others point toward alternative ways to improve reliability and user experience. Table 6 shows how the number of GenAI exchanges increases with a user’s skill, and users tend to prefer Iterative conversations to Comprehensive ones (Section 3.2.4). Accordingly, high-profile tools usually facilitate iteration in some respect, most obviously by allowing users to send multiple prompts throughout a conversation. However, there are other ways to facilitate iteration, such as allowing users to edit earlier prompts to create conversation branches.
Our findings indicate that daily use makes productivity improvements significantly more likely (Section 3.1.2), and also correlates with GenAI use on a broader range of tasks. This suggests that the best way for individual developers to realize the benefits of these new tools is to carefully consider which tasks in their workflow could benefit from utilizing GenAI, then purposefully learn about and apply GenAI there. Gradual adjustment may be necessary to ensure consistent performance and build an understanding of how best to navigate AI limitations.
4.3. High Technique Awareness, Low Adoption
Our survey population had an overwhelming familiarity with various prompting techniques, with their awareness of the Meta Language Creation technique being the lowest (74%). However, despite an overall high awareness of prompting techniques, there was a lackluster adoption. For example, even though 92% (66/72) of respondents indicated familiarity with the Output Style technique, only 82% of those (54) actually used it at all, and of those only 54% (29) used it regularly. The adoption percentages are worse for the remaining techniques, and regular usages are well below 50%. This is particularly interesting given that awareness of research outcomes is typically one of the biggest hurdles for tool and technique adoption, meaning that the issue must lie elsewhere. It may be that either developers do not find these techniques helpful or that they are unaware of how to effectively use them. The results in Table 4 suggest that decades of work experience may assist veteran developers in finding use cases for prompting techniques, since this was the group most likely to adopt these techniques. However, flexibility may also be a crucial component to effective use, as potentially demonstrated by the intermediate developers who had the highest frequency of regular technique usage. Accordingly, future research can further explore the pain points that developers encounter when applying these techniques to their workflows. Additionally, a comprehensive taxonomy and/or usage guide could potentially assist developers in more effectively using these prompting techniques or understanding how to apply them to their particular use cases.
4.4. Future Work
While existing work provided a helpful starting point for the prompting techniques included in the survey, it was not comprehensive enough to account for all interactions, necessitating our definitions of Output Style and Condition Check. Despite these additions, our qualitative analysis revealed strategies not adequately represented in the survey. One user instructed the GenAI to avoid outputting explanations of its work, while another deliberately excluded examples of input and output from prompts, suggesting that some users develop preferences for concise interactions or that extra context can reduce model output quality in certain situations.
To rigorously study the effectiveness, variation, and progression of GenAI interactions, a comprehensive taxonomy of all possible interaction types must be created. This would enable systematic categorization and quantification of strategies, allowing controlled comparisons of effectiveness across tasks and user populations.
Given iteration’s significant correlation with skill and productivity, investigating further iterative interaction methods could be valuable. For example, allowing users to edit model parameters (temperature, output length, etc.) mid-conversation could open new solution paths and grant greater control of GenAI tools.
5. Threats to Validity
Internal Validity. To mitigate bias, we employed an iterative open-coding methodology for analyzing open-ended survey responses. We also followed best practices when formulating our survey questions in order to avoid confusing and/or biasing language, and conducted a small pilot study with graduate students to test the content, clarity, and length of our questionnaire. We recruited participants from a variety of sources (e.g., our professional networks, university organizations, etc.) to ensure we received a broad range of perspectives and experiences, but we are aware of the threat of self-selection bias.
External Validity. The conclusions of this study only apply to those participants that completed our survey and cannot be overly generalized to the larger developer population. However, given the themes and trends that emerged, we believe that many developers will share similar experiences to those reported. Ultimately, our goal was never to claim generalizability, but instead to explore the emerging landscape of how developers are interacting with GenAI tools for SE tasks.
Construct Validity. This work relies on self-reported developers’ perceptions, which can be subjective and influenced by recall and social desirability biases. We sought to mitigate this threat by corroborating and anchoring subjective assessments such as proficiency with more concrete metrics captured within the survey, including the breadth of AI-assisted tasks participants perform and their frequency of tool use, and avoiding making definite claims about practices. The demographic information provided in our study was supplied by respondents and cannot be independently verified.
6. Related Work
Generative AI models are systems capable of creating new content based on their training data. Although the definitions of ‘prompting’ vary, they generally refer to providing GenAI with input, usually text instructions and / or an image, to guide the model toward producing the intended result. Upon release, these tools were quickly applied to SE, with an early study by GitHub on their Copilot integration (Eirini Kalliamvakou, [n. d.]) showing significant gains in coding speed, job satisfaction, and perceived productivity. A more recent mixed-methods study from Butler et al. (Butler et al., 2024) within a large software company showed an increase in positive views toward AI coding tools, but without a corresponding boost to developer trust, which is notable because trust factors play an important role in developer intentions (Choudhuri et al., 2024) (Choudhuri et al., 2025b). The utility of AI tools goes beyond coding tasks, as APIs or locally run models have led to a new design paradigm of integrated, interacting commands with more traditional structured code (Liang et al., 2025; Tafreshipour et al., 2025). Multiple reviews of LLM (Hou et al., 2024; Fan et al., 2023) and LLM-based agents (Liu et al., 2024) for SE show extensive possibilities to further improve SE workflows.
This has created the field of prompt engineering, which serves to determine the most effective ways of prompting a model in order to produce high-quality output which suits the user’s needs. In its early stages, this resulted in explorations of individual strategies, leading to catalogs of prompt patterns (White et al., 2023) applicable to a broad range of interaction types (Treude and Gerosa, 2025). As more patterns were refined, improved and discovered, compilations of techniques were formed through extensive reviews of the academic literature (Schulhoff et al., 2025; Sahoo et al., 2025), though empirical analysis has found no significant differences in code quality across basic prompt patterns (Della Porta et al., 2025). Others attempted to apply prompt engineering for specific applications such as security (Tony et al., 2025), design more complex conversational workflows (Robino, 2025), and even use the generative AI models themselves for prompt engineering tasks (Zhou et al., 2023). With new advances such as chain-of-thought outputs (Treude and Kula, 2025), human-AI interaction is evolving beyond just prompting.
Although these frameworks and pattern collections have provided an important foundation for understanding prompting capabilities, they reveal little about how software developers use AI tools in practice. The DevGPT dataset (Xiao et al., 2024) and various analyses of it have sought to close this gap, but often focus on a broader analysis of collaborative contexts (Hao et al., 2024), specific tasks such as issue resolution (Ehsani et al., 2025), or high-level usage patterns such as autocomplete / coding acceleration (Liang et al., 2024) rather than an examination encompassing diverse types of SE tasks. Li et al. conducted a comprehensive study on the individual and organizational motives of AI adoption and use among software developers (Li et al., 2024), while Simaremare et al. highlighted the key benefits and challenges of doing so (Simaremare and Edison, 2024). Other approaches involve analysis of discourse from social media sites such as X/Twitter (Basha and Rodríguez-Pérez, 2025) or user reviews about AI coding extensions (Lyu et al., 2025), while surveys of AI adoption among developers (Russo, 2024; Lambiase et al., 2025; Zakharov et al., 2025) in Sweden (Lakis and Rifai, 2025), Slovenia (Robas, 2025), and Indonesia (Anditama and Hidayanto, 2024) reveal factors influencing use and perceptions among certain populations.
While many distinct techniques have been researched and compiled, there has been no investigation comparing developers’ preferred prompting strategies and interaction methods across multiple categories of software engineering tasks, leaving information about how these tools are used in practice relatively scarce. Our study addresses this gap with a comprehensive analysis of prompting and conversation practices in six distinct SE domains.
7. Conclusion
We conducted a comprehensive survey of 72 software developers to investigate how they integrate GenAI tools into software engineering workflows. Our findings reveal primary adoption of GenAI for code generation, while use for related tasks such as documentation and debugging strongly correlates with a developer’s AI proficiency and usage frequency. Users tend to prefer iterative conversations to achieve their goals, with consistent single-shot prompting remaining an impossibility. Significant reliability challenges remain for complex coding tasks, particularly debugging.